# AfriDoc-MT: Document-level MT Corpus for African Languages

### v1

@@ 4.2 Experimental Setup
Given that our created dataset can be used for sentence-level translation and as a baseline for document-level translation, we evaluate all models on the test splits for each domain. We evaluate the translation models (M2M-100, NLLB-200, and MADLAD-400) using the Fairseq Ott et al. (2019) codebase for (M2M-100 and NLLB-200), and the Transformers Wolf et al. (2020) codebase for MADLAD-400. However, for other models including Aya-101, we use the EleutherAI LM Evaluation Harness (lm-eval) tool (Biderman et al., 2024) using the three templates listed in Table 19 of Section B.4.
@@
Going further, we conduct document-level translation using a similar setup as the sentence-level experiment, but on a few selected models as ideally not all the models have the context length requirement to handle the translation of entire documents. An initial analysis revealed that some models were unable to process entire documents due to their maximum input length being shorter than the token counts for some languages, particularly African languages such as Amharic and Yorùbá. To address this, we adopted a pseudo-document approach, splitting documents into smaller, fixed-size chunks of sentences that fit within the models’ token limits. The final chunk in the document could contain fewer than sentences. We experimented with different chunk sizes ( = 5, 10, 25), with =1 serving as the sentence-level setup. Based on our findings, we selected =10 and used this setup for our experiments unless stated otherwise. Table 10 shows the resulting number of parallel pseudo-documents and the average and 95th percentile token counts per pseudo-document for each language and model tokenizer.
@@ 4.3 Evaluation Metrics
The evaluation of document-level translation remains a challenge due to the inability of existing automatic metrics to indicate document-level improvements and identify discourse phenomena Jiang et al. (2022); Dahan et al. (2024), while embedding-based metrics have not yet been explored in the context of African languages. We computed our document-level metrics by first realigning either sentence-level or pseudo-translation outputs into complete documents. Then, we applied the vanilla BLEU and CHRF metrics to these realigned documents, which we refer to as document BLEU (d-BLEU) Papineni et al. (2002) and document CHRF (d-CHRF) Popović (2015). The metrics are calculated using SacreBLEU999case:mixed|eff:no| tok:13a|smooth:exp|v:2.3.1, Post (2018) with significance tests based on bootstrap resampling, reporting the confidence intervals for the scores from a sample size of . We report the d-CHRF scores for the best prompt for each model and language direction in the main text, with all additional results provided in the Appendix C.
@@
Furthermore, we use GPT-4o as a proxy for human evaluation to evaluate the translation outputs. Recent work has demonstrated that LLMs can effectively assess translation quality and provide analyses of translation errors Wu et al. (2024); Sun et al. (2024). Following a similar approach to Sun et al. (2024), we use GPT-4o to assess Fluency, Content Errors (CE), and Cohesion Errors—specifically lexical (LE) and grammatical (GE) errors. However, due to cost constraints, we limit this evaluation to a few model outputs. We provide more details in Section B.6.
@@ 5.1 Sentence-level Evaluation
In Tables 5 and 4 we present d-CHRF scores based on the realigned documents, created by merging the translated sentences into their corresponding documents. We highlight our main findings below, and sentence-level evaluation results using sentence-level metrics are reported in Appendix C.
| Model | Setup | eng X | X eng | ||||||||
| d-CHRF | Fluency | CE | LE | GE | d-CHRF | Fluency | CE | LE | GE | ||
| Aya-101 | Sent | ||||||||||
| Doc10 | |||||||||||
| GPT-3.5 | Sent | ||||||||||
| Doc10 | |||||||||||
| LLaMAX3-SFT1 | Sent | ||||||||||
| Doc10 | |||||||||||
| LLaMAX3-SFT10 | Doc10 | ||||||||||
| Sent | |||||||||||
| GPT-4o | Doc10 | ||||||||||
@@
On average the NLLB models obtain scores of 65.4/66.6 and 64.3/65.0 on health and tech domains respectively, with 3.3B outperforming 1.3B except when translating into Yorùbá. When translating to English, the worst performing model across the two domains is Toucan, however, translating to African languages gives better results than MADLAD-400 and Aya-101. Furthermore, translating to African languages is significantly worse compared to translating to English for all the models.
@@ 5.2 Document-level Evaluation
In Tables 7 and 6 we present d-CHRF scores based on the best prompt per language for the the translation output of the models when evaluated on the realigned documents from pseudo-documents with 10 sentences per pseudo-document.
@@
Both LLama models trained via SFT on sentences (LLama3.1-SFT, and LLaMAX3-SFT) show a decline in performance in the pseudo-document setting compared to sentence-level translation. However, the same models trained via SFT on pseudo-documents with =10 demonstrate significant improvements on pseudo-documents. Interestingly, the LLaMAX3-SFT10 performs consistently well, achieving results comparable to its sentence-level counterpart on sentence-level tasks, and also outperforming LLama3.1-SFT10, particularly when translating into African languages.
@@ 5.3 GPT-4o based evaluation
Table 8 presents the results of the GPT-4o evaluation of realigned documents from both sentence-level tasks and pseudo-document-level tasks (with k=10), focusing on outputs from selected models and translations between English and Zulu/Swahili in the health domain. Our findings indicate that, as anticipated, GPT-4o, when acting as the evaluator, consistently rates its own outputs as the best across all metrics, suggesting potential self-bias. It rates the pseudo-document outputs as more fluent than the sentence-level outputs for both translation directions for all the models. Similarly, the result shows that translating pseudo-documents results in fewer content errors, lexical errors, and grammatical errors. Lastly, LLaMAX3-SFT10 has the best evaluation results even in cases where d-CHRF does not rate it as best. We discuss this further in Section C.3.
@@ 6 Discussion and Analysis
In order to better understand the models’ behaviour, we analyse their outputs based on frequently observed problems in document-level MT using LLMs Wu et al. (2024); Wang et al. (2024b). We conduct these analyses on the models’ pseudo-document (10) translation outputs before merging them into their actual documents, unless stated otherwise. We provide more results in Appendix D.
@@
We analyzed the translation outputs comparing them to their corresponding reference translation to check wether they are empty or if the models under-generate. Our analysis shows that all models rarely generate empty translations (refer to Appendix D). However, GPT-3.5 and GPT-4o exhibit a slight tendency to generate empty translations when translating into Yorùbá and Zulu for both domains, though this occurs rarely (less than 10%). For output length, translations with lengths less than 70% of the reference translations were considered under-generated. Figure 1 shows that all the models have the tendency to under-generate. Aya-101 under-generates more than 3% of the pseudo-documents when translating into all the languages from English. LLaMAX3-Alpaca shows at least 10% under-generation across the languages, while other models have less than 30% when translating into English.
@@
We evaluate whether these models understand the task by generating outputs in the target languages using a language identification task. Our results show that these models rarely generate outputs in the wrong language when translating to English. However, when translating to African languages, there is a higher likelihood of incorrect language translations, particularly with open models (Figure 2).
@@
We compare the average d-CHRF scores obtained by selected models, including GPT-3.5/4 and LLama3.1-SFTk where 1,5,10. The evaluation was conducted across all pseudo-document lengths: 1, 5, 10, 25, and the full length. Figure 3 shows that for translations into African languages, d-CHRF scores decrease as document length increases. A similar trend is observed for the reverse translation, except for GPT-4o, which shows an increasing trend. Also, models trained on long documents generalize better to long documents than those trained on sentences.
What language benefits more from supervised finetuning?
We focus on the sentence-level task and translated across all 30 directions for which the model was trained, evaluating both NLLB-200 (1.3B) and its fine-tuned version using d-CHRF. Figure 4 shows performance improvements after supervised fine-tuning of NLLB-200 for both domains. The results shows that translating into Yorùbá, which is the direction with the lowest d-CHRF score from English among all the languages, benefited the most. One major factor contributing to this is the presence of diacritics.
@@ 7 Conclusion
In this work, we present AfriDoc-MT, a document-level translation dataset covering two domains health and tech for 5 African languages. We conducted document-level translation benchmarks, evaluating models of various sizes and fine-tuning selected ones. Due to context length limitations, documents were translated in two ways: (1) sentence by sentence and (2) as pseudo-documents. Outputs were assessed using classical MT metrics and GPT-4o as a proxy for human evaluation. Among built-in MT models, NLLB-200 showed the best performance, while GPT-4o outperformed general-purpose LLMs, with fine-tuning of selected models yielding significant improvements. GPT-4o found pseudo-document translations more fluent and accurate than sentence-level ones, contrasting with classical metrics and highlighting the need for better evaluation metrics. Our analysis also reveals that some LLMs are prone to generating short outputs, off-target translations, and that languages do not benefit equally from fine-tuning.
@@ 8 Limitations
We evaluated only a small subset of the numerous multilingual LLMs available. Our experiments were also limited by the context length of the LLMs, particularly for open LLMs. Except for LLama3.1, all other open LLMs have a context length of 8192 tokens, while encoder-decoder models were primarily based on T5. This makes it difficult to use the context length beyond a certain limit, making full document translation infeasible. Additionally, LLMs are prone to variance in performance based on the prompt; therefore, we evaluated them for translation using three different prompts. However, it is possible that our prompts were not optimal.
@@
Africa is home to thousands of indigenous languages, many of which exhibit unique linguistic properties. However, due to the high cost of translation using human translators and limited available funding, it is currently impossible to cover all languages. As a result, we focused on just five languages. We hope that future work will expand this dataset to include more languages and inspire the creation of additional datasets with broader coverage for document-level translation. Similarly, AfriDoc-MT is a multi-way parallel dataset. However, due to the cost of running inference over three prompts and across all 30 translation directions for all the models evaluated, most of our analysis is limited to translation tasks between English and the five African languages. While we fine-tuned NLLB-200, LLama3.1 and LLaMAX3 on all 30 directions, we only provide results from NLLB-200 for all directions both before and after fine-tuning for sentence-level and pseudo-document tasks in the Appendix.
@@
Quality evaluation in machine translation is an open and ongoing area of research, especially for document-level translation. Recent works have proposed embedding-based metrics for evaluation at both the sentence and document levels. While this has been well explored for high-resource language pairs, it remains underexplored for African languages, although there is a tool, AfriCOMET, that works for sentence-level evaluation in African languages. Furthermore, we did not carry out human evaluation due to the cost; instead, we used GPT-4 as a proxy. The model’s understanding of these languages is not well established, although it achieves comparably the best performance when compared to other decoder-only LLMs. Furthermore, we evaluated a few models and only 3 languages using GPT-4o due to resource constraint.

### v2

@@ 4.2 Experimental Setup
Given that our created dataset can be used for sentence-level translation and as a baseline for document-level translation, we evaluate all models on the test splits for each domain. We evaluate the translation models (M2M-100, NLLB-200, and MADLAD-400) using the Fairseq (Ott et al., 2019) codebase for (M2M-100 and NLLB-200), and the Transformers (Wolf et al., 2020) codebase for MADLAD-400. However, for other models including Aya-101, we use the EleutherAI LM Evaluation Harness (lm-eval) tool (Biderman et al., 2024) using the three templates listed in Table˜23 of Section˜B.4.
@@
We also perform document-level translation using a setup similar to the sentence-level experiment, but only with models that meet context length requirements. An initial analysis showed that some models were unable to process entire documents due to input length limits, which were exceeded by token counts in some languages (Amharic and Yorùbá). To address this, we adopted a similar approach to Lee et al. (2022), splitting documents into fixed-size chunks of sentences to fit within token limits; the final chunk may contain fewer than sentences. To select an appropriate chunk size, we conducted initial tests with (sentence-level), 5, 10, and 25, choosing for our experiments. We provide results from this analysis in Table˜11.
@@ 4.3 Evaluation Metrics
Evaluating document-level translation remains challenging, as existing automatic metrics struggle to capture improvements and account for discourse phenomena (Jiang et al., 2022; Dahan et al., 2024), and embedding-based metrics have not been explored in this context for African languages due to the lack of data. Hence, we realigned sentence-level or pseudo-translation outputs into full documents, then computed BLEU (Papineni et al., 2002) and chrF (Popović, 2015) to create document BLEU (d-BLEU) and document chrF (d-chrF). Metrics were computed using SacreBLEU101010case:mixed|eff:no|tok:13a|smooth:exp|v:2.3.1 (Post, 2018) with bootstrap resampling () to report 95% confidence intervals. We report d-chrF scores for the best prompt per model and language direction in the main text, as chrF better captures the morphological richness of African languages (Adelani et al., 2022), with full results provided in Appendix˜C.
@@
We use GPT-4o as a judge to evaluate translation outputs, following recent work showing LLMs’ effectiveness in assessing translation quality and analyzing errors (Wu et al., 2024; Sun et al., 2025). Following Sun et al. (2025), we assess each translated document’s fluency, content errors (CE), and cohesion errors—specifically lexical (LE) and grammatical (GE) errors—using GPT-4o, with evaluation limited to a few model outputs due to cost constraints (Section˜B.6). We also complement this with human evaluation for direct assessment scores (Section˜B.7) and qualitative analysis through manual inspection (Section˜B.8).
@@ 5.1 Sentence-level Evaluation
In Tables˜6 and 5 we present d-chrF scores based on the realigned documents, created by merging the translated sentences into their corresponding documents. We highlight our main findings below, and sentence-level evaluation results using sentence-level metrics are reported in Appendix˜C.
| Model | Setup | eng X | X eng | ||||||||||
| amh | hau | swh | yor | zul | avg | amh | hau | swh | yor | zul | avg | ||
| Fluency | |||||||||||||
| Aya-101 | Sent | 2.8 | 2.9 | 1.3 | 1.2 | 2.7 | 2.2 | 3.0 | 2.9 | 3.3 | 2.3 | 2.9 | 2.9 |
| Doc10 | 2.5 | 2.9 | 1.8 | 1.4 | 3.1 | 2.3 | 3.6 | 3.2 | 3.8 | 2.9 | 3.3 | 3.4 | |
| GPT-3.5 | Sent | 1.1 | 1.4 | 4.9 | 1.1 | 1.5 | 2.0 | 3.0 | 2.5 | 3.6 | 2.6 | 2.6 | 2.9 |
| Doc10 | 1.0 | 1.1 | 4.9 | 1.0 | 1.3 | 1.9 | 4.2 | 4.1 | 4.7 | 3.9 | 4.0 | 4.2 | |
| LLaMAX3-SFT1 | Sent | 3.5 | 3.2 | 3.6 | 4.0 | 3.3 | 3.5 | 3.5 | 2.9 | 3.9 | 3.2 | 3.6 | 3.4 |
| Doc10 | 1.7 | 2.3 | 3.2 | 2.7 | 2.4 | 2.5 | 3.0 | 2.6 | 3.1 | 3.2 | 3.1 | 3.0 | |
| LLaMAX3-SFT10 | Doc10 | 2.6 | 3.9 | 4.4 | 4.2 | 3.6 | 3.8 | 4.1 | 4.2 | 4.6 | 4.4 | 4.4 | 4.3 |
| Content Error | |||||||||||||
| Aya-101 | Sent | 9.0 | 9.1 | 9.5 | 8.2 | 13.7 | 9.9 | 15.8 | 17.1 | 17.5 | 23.8 | 19.1 | 18.7 |
| Doc10 | 8.9 | 8.3 | 8.2 | 6.7 | 16.1 | 9.6 | 12.6 | 14.6 | 14.1 | 18.7 | 15.7 | 15.1 | |
| GPT-3.5 | Sent | 7.2 | 11.3 | 4.2 | 7.8 | 20.9 | 10.3 | 9.5 | 13.2 | 12.4 | 12.9 | 15.9 | 12.8 |
| Doc10 | 3.3 | 7.4 | 3.9 | 4.1 | 10.1 | 5.8 | 6.6 | 9.2 | 7.1 | 9.0 | 10.8 | 8.5 | |
| LLaMAX3-SFT1 | Sent | 10.0 | 9.5 | 12.3 | 12.3 | 12.4 | 11.3 | 11.5 | 9.6 | 11.8 | 12.2 | 12.5 | 11.5 |
| Doc10 | 10.4 | 8.8 | 9.5 | 8.1 | 8.4 | 9.0 | 9.0 | 9.0 | 8.9 | 9.1 | 8.6 | 8.9 | |
| LLaMAX3-SFT10 | Doc10 | 7.3 | 14.6 | 10.5 | 10.3 | 12.1 | 11.0 | 8.7 | 9.6 | 8.9 | 10.2 | 9.4 | 9.4 |
@@
On average the NLLB models obtain scores of 65.4/66.6 and 64.3/65.0 on health and tech domains respectively, with 3.3B outperforming 1.3B except when translating into Yorùbá. When translating to English, the worst performing model across the two domains is Toucan. However, it gives better results than MADLAD-400 and Aya-101 when translating to African languages. Furthermore, translating to African languages is significantly worse compared to translating to English for all the models.
@@ 5.2 Document-level Evaluation
In Tables˜8 and 7 we present d-chrF scores based on the best prompt per language for the translation output of the models when evaluated on the realigned documents from pseudo-documents with 10 sentences per pseudo-document.
@@
Both LLama models trained via SFT on sentences (LLama3.1-SFT, and LLaMAX3-SFT) show a decline in performance in the pseudo-document setting compared to sentence-level translation. However, the same models trained via SFT on pseudo-documents with =10 demonstrate significant improvements on pseudo-documents. Interestingly, the LLaMAX3-SFT10 model performs consistently well, achieving results comparable to its sentence-level counterpart on sentence-level tasks, and also outperforming LLama3.1-SFT10, particularly when translating into African languages.
@@ 5.3 GPT-4o based evaluation
Table˜9 presents average GPT-4o evaluations for fluency and content errors (CE) of realigned outputs from sentence-level and pseudo-document-level tasks (k=10) across four models in the health domain. When translating into English, pseudo-document outputs are generally rated as more fluent and show fewer content errors, except for LLaMAX3-SFT1, which, when evaluated on pseudo-documents, shows lower fluency but still fewer content errors—an outcome that is counterintuitive. However, when translating into African languages, the results are less consistent. Notably, GPT-3.5 achieves a fluency score of 4.9 for both fine-tuned versions of LLaMAX3-SFT—a score no model achieved when translating into English. Additionally, Yorùbá, a language that had some of the lowest d-chrF scores across models, achieved fluency scores of 4.0 and 4.2 with the two LLaMAX3-SFT versions. These inconsistencies raise concerns about GPT-4o’s reliability. Consequently, we focus on human evaluation going forward. Full GPT-4o results are provided in Section˜C.3.
5.4 Human evaluation
In Table˜10 we report average direct assessment (DA) scores (on a scale from 0 to 100) from three annotators per language for the health domain, when translating into four African languages. For each language, we used 30 documents across models and both domains to compute inter-annotator agreement. We obtained Krippendorff’s alpha values of 0.40, which are relatively low due to the fine granularity of the evaluation scale. Human evaluation results align closely with d-chrF, which favors sentence-level translations over pseudo-document translations when translating into African languages. Among the models, LLaMAX3-SFT1 receives higher ratings at the sentence-level but is rated lower when translating pseudo-documents. In contrast, LLaMAX3-SFT10 receives slightly lower ratings than LLaMAX3-SFT1 at the sentence-level but is rated higher in the pseudo-document setting. GPT-3.5 is generally rated as the weakest model across languages, except for Swahili, where its performance is comparatively better (see Section˜C.4 for details).
| Model | Setup | amh | hau | swh | yor | zul |
| GPT-3.5 | Sent | 14.6 | 29.6 | 66.5 | 7.5 | 9.2 |
| Doc10 | 1.7 | 16.4 | 68.3 | 4.2 | 3.1 | |
| LLaMAX3-SFT1 | Sent | 64.5 | 81.5 | 57.9 | 65.1 | 48.1 |
| Doc10 | 27.4 | 45.7 | 50.6 | 44.3 | 22.5 | |
| LLaMAX3-SFT10 | Doc10 | 38.5 | 76.7 | 62.4 | 64.9 | 46.7 |
5.5 Qualitative evaluation
Our qualitative analysis, based on feedback from native speakers who are also authors, indicates that GPT-3.5 frequently over-generates in the pseudo-document setup by repeating words and phrases—except in Swahili, where it performs best. However, for Yorùbá, it often uses inconsistent or partial diacritics, resulting in inaccuracies. LLaMAX3-SFT1 also exhibits repetition in pseudo-document translations, likely due to a length generalization problem (Anil et al., 2022), and does so more than LLaMAX3-SFT10. For the other four languages, LLaMAX3-SFT1 with the sentence-level setup was rated higher than other models and configurations, owing to better context preservation and fewer repetitions. These observations are consistent with both d-chrF and DA scores, although d-chrF scores tend to be inflated.
@@ 6 Discussion and Analysis
To better understand model behavior, we analyze their pseudo-document () translation outputs based on our findings and common issues in document-level MT with LLMs (Wu et al., 2024; Wang et al., 2024b). Additional results are provided in Appendix˜D.
@@
We compare model translations to the reference translations to identify empty outputs and cases of under- or over-generation. We found that all models rarely generate empty translations (refer to Appendix˜D), although GPT-3.5 and GPT-4o showed a slight tendency to do so for Yorùbá and Zulu, occurring in under 10% of cases. To evaluate this we compute the compute model output to reference translation length ratio across five models and for the languages, for the evaluated models and took the lowest 10th percentile, and this returns 73% and upper percentile of 550%. However, we defined under-generation as outputs <70% of reference length, and over-generation as >143%. As shown in Figures˜1(a) and 1(b) and consistent with our qualitative findings, GPT-3.5 tends to over-generate more in African languages except for Swahili, while LLaMAX3-SFT1 often under-generates as a sentence-level model. Moreover, all models over-generated by about 20% for Amharic, likely due to its unique script.
@@
We evaluate whether these models understand the task by generating outputs in the target languages using the OpenLID (Burchell et al., 2023) language identification model. Our results show that these models rarely generate outputs in the wrong language when translating to English. However, when translating to African languages, there is a higher likelihood of incorrect language translations, particularly with open models (see Figures˜1(c) and 1(d)). These off-target languages often include English, and other languages including other African languages.
@@
We compare average d-chrF scores of selected models, including GPT-3.5/4 and LLama3.1-SFTk ( = 1, 5, 10), evaluated across pseudo-document lengths of 1, 5, 10, 25, and full length. As shown in Figure˜2, d-chrF scores generally decline with increasing document length for African language translations. The reverse translation direction shows a similar trend, except for GPT-4o, which improves with length. Models trained on longer documents also generalize better to longer inputs than those trained on sentences.
@@ 7 Conclusion
We introduce AfriDoc-MT, a document-level translation dataset in the health and tech domains for five African languages. We benchmarked various models, fine-tuning selected ones. Due to context length limits, documents were translated either sentence by sentence or as pseudo-documents. Outputs were evaluated using standard MT metrics, GPT-4o as a judge, and human direct assessment. NLLB-200 was the strongest built-in MT model, while GPT-4o outperformed general-purpose LLMs. However, our DA and qualitative analysis found GPT-4o’s judgments inconsistent for African languages, and sentence-by-sentence translation proved more effective for some languages.
@@ 8 Limitations
We evaluated only a small subset of the numerous multilingual LLMs available. Our experiments were also limited by the context length of the LLMs, particularly for open LLMs. Except for LLama3.1, all other open LLMs have a context length of 8192 tokens, while encoder-decoder models were primarily based on T5. This makes it difficult to use the context length beyond a certain limit, making full document translation infeasible. Additionally, LLMs are prone to variance in performance based on the prompt. We therefore evaluated them for translation using three different prompts. However, it is possible that our prompts were not optimal.
@@
Africa is home to thousands of indigenous languages, many of which exhibit unique linguistic properties. However, due to the high cost of translation using human translators and limited available funding, it is currently impossible to cover all languages. As a result, we focused on just five languages. We hope that future work will expand this dataset to include more languages and inspire the creation of additional datasets with broader coverage for document-level translation. Similarly, AfriDoc-MT is a multi-way parallel dataset. However, due to the cost of running inference over three prompts and across all 30 translation directions for all the models evaluated, most of our analysis is limited to translation tasks between English and the five African languages. While we fine-tuned NLLB-200, LLama3.1 and LLaMAX3 on all 30 directions, we only provide results from NLLB-200 for all directions both before and after fine-tuning for sentence-level and pseudo-document tasks in the Appendix˜D.
@@
Quality evaluation in MT is an open and ongoing area of research, especially for document-level translation. Recent works have proposed embedding-based metrics for evaluation at both the sentence and document levels. While this has been well explored for high-resource language pairs, it remains underexplored for African languages, although there is a tool, AfriCOMET, that works for sentence-level evaluation in African languages. However, we evaluated the document-level translation outputs using ModernBERT-base-long-context-qe-v1111111https://huggingface.co/ymoslem/ModernBERT-base-long-context-qe-v1, trained on the WMT human evaluation dataset across 41 language pairs, including over 20 languages and three African languages (Hausa, Xhosa, and Zulu), two of which are included to our work. However, the scores were nearly identical across all models, offering no meaningful differentiation. Hence, for our document-level evaluation, in addition to lexical-based metrics, we incorporated three other evaluation approaches: using GPT-4o as a judge, human evaluation, and qualitative analysis. GPT-4o was employed to assess and rate the translation outputs of four models. While its ratings were consistent for translations into English, the same was not observed for translations into African languages, likely due to the model’s limited understanding of these languages. Therefore, we conducted a human evaluation for translations from English to African languages, comparing only three models due to cost constraints. However, we were unable to recruit annotators for Zulu.
Model Coverage and Evaluation Focus
While we fine-tuned both NLLB-1.3B and LLaMAX3 models across all 30 language directions, due to computational constraints and the high cost of qualitative evaluation, our detailed analysis focuses only on translation between English and the 5 African languages. Nevertheless, we report quantitative results across all 30 directions for NLLB-1.3B. We will make all fine-tuned models publicly available to support future work, and we hope that further research will explore the remaining translation directions in greater depth.
Translationese and English-Centric Bias
A potential limitation of our dataset is the influence of translationese (Koppel and Ordan, 2011). Since all source material translated originates in English, translated sentences in African languages may exhibit patterns such as unnatural syntax or overly literal phrasing. Although we have not conducted an analysis to quantify these effects, prior work suggests that they can affect MT model performance, generalization and evaluation including direct assessment (Freitag et al., 2019; Edunov et al., 2020). Furthermore, AfriDoc-MT may reflect a bias toward English in terms of structure, semantics, and cultural framing. We leave a deeper investigation of these issues to future work.